OpenStudy (anonymous):

solve the proportion 5/3c=1/6

OpenStudy (anonymous):

solve for c?

OpenStudy (anonymous):

if so, then it is 1/10=c

OpenStudy (australopithecus):

5/3c = 1/6 5 = 3c(1)/6 5*6 = 3c(1) 30 = 3c 30/3 = c 10 = c

OpenStudy (anonymous):

ahh, my bad